# thanks: – corresponding author

### v2

@@ 1 Introduction
The presented description summarizes the main results of the one-particle picture of the squeezed coherent state construction. Applications and the role played by coherent states in quantum optics are described in Ref. [12, 13]. Other applications of the coherent states cover wide range of mathematical physics issues, starting from quantization [2, 14], through cosmology [15, 16], to field theory [17].
Another look at the coherence states and their properties comes from the many-particle perspective. However, this exciting problem states a different thread we do not develop here because it is out of the scope of the proposed research. So, instead of this, we focus on the simplest case; namely, we consider the two-particle problem. It is admittedly more straightforward but still offers the possibility of a deeper insight into the entanglement phenomena and the role of quantum coherence, noted as intriguing and unsolved problem in Refs. [18, 19]. The structure of the bipartite states realized with the help of the squeezed coherent states opens a way to study the quantum states, containing entanglement and coherence on an equal footing. In other words, these two phenomena coexist in such a state through its construction.
@@
The novelty of the presented approach is that it examines the squeezing and entanglement phenomena starting from the description of bipartite states in the Bargmann space. The first step enables us to establish a direct connection between the parameter, $\alpha$, associated with the non-rotational measure in the Bargmann space and the squeezing parameter. By crossing with the calculations to the phase space and applying the results of Simon and de Gosson, we associate the phenomena of squeezing and entanglement. Moreover, we quantify the degree of entanglement using the log-negativity criterion using the symplectic eigenvalues [37] of the covariance matrix modified by Simon [38]. Our result explicitly depends on the parameter $\alpha$, thereby affirming the connection between the parameter associated with the non-rotational measure in the Bargmann space and entanglement. Finally, we demonstrate that the system under investigation can be represented by two quantum-mechanical coupled anisotropic harmonic oscillators, with the coupling and anisotropy originating from the squeezing parameter. This result complements study on the application of the coupled harmonic oscillators to investigate the problem between entanglement and squeezing in continuous-variable Gaussian systems [39]. This one simultaneously refers to the examination of the two coupled harmonic oscillators in the Segal-Bargmann space representation presented in Ref. [40].
The rest of this paper is organized as follows. In Sec. 2, we recall selected properties of the holomorphic Hermite functions and discuss their utility in constructing single- and bipartite coherent and squeezed states. The main results of our study and their discussion are presented in Sec. 3. In the first part of this section, we use the phase-space approach based on the Wigner distribution function to quantitatively investigate the entanglement issue in continuous-variable systems. In turn, in the second part, we develop a theoretical model of a two-particle mechanical system illustrating details of our study. Section 4 contains a summary and conclusions. Finally, the paper concludes with an appendix containing the proofs of the Mehler formulae.
2 Mathematical preliminaries
@@
[…]
@@
[…]
@@
The second class of analytic functions, $\varphi_{2,(m,n)}^{(\alpha)}(z_{1},z_{2})$, to be considered by us is generated by the holomorphic Hermite functions of the two complex variables, $h_{m,n}^{(\alpha)}(z_{1},z_{2})$, with $m,n=0,1,\ldots$, and $10<\alpha<$ 1, i.e. we assume that
@@
whose orthogonality is treated by [34, Theorem 3]. We emphasize that the Hermite functions (22) form the orthonormal basis with respect to the Gaussian measure (12). The algebraic and analytic properties of these polynomials are considered in Refs. [34, 43, 44, 45, 46]. In particular, it is also worth noting that the complex Hermite polynomials, $H_{m,n}(z_{1},z_{2})$ for $z_{2}=\bar{z}_{1}$ reduce themselves to the polynomials of the real variables $x,y$ with complex coefficient [42, 47, 48]. For the first time, this kind of polynomials, i.e. $H_{m,n}(z_{1},{\bar{z}_{1}})$ appeared in Ito’s work [47] on applications of complex random variables in stochastic processes. Hence, they are called the Ito polynomials. Finally, note that both the functions given by Eq. (17) and Eq.(21) satisfy the condition in the form
$\sum_{m,n=0}|\varphi_{k,(m,n)}^{(\alpha)}(z_{1},z_{2})|^{2}<\infty,\quad\textrm{for all $z_{1},z_{2}\in\mathbb{C}$},$ (24)
@@
3 Results and discussion
3.1 Wigner representation
The phase-space representation is often used to examine physical systems, such as electromagnetic modes in optical or microwave cavities, or mechanical modes of harmonic oscillators, where degrees of freedom constitute continuous variables. This approach, frequently applied in quantum optics, is intuitive and robust, facilitating the investigation of the system non-classical state features or their dynamics. According to this approach, the system states within the phase space are represented by non-classical distribution functions of conjugate variables defined on a real symplectic space $\mathbb{R}^{2}\times\mathbb{R}^{2}$ equipped with cannonical symplectic form, relevant to these variables. Among numerous such functions, the Wigner distribution function (WDF) plays a pivotal role, providing a lucid phase-space representation of the quantum state and emphasising the cognitive aspect of the obtained results. Let us first start by introducing the WDF for a bipartite pure state $\left|f\right>$ represented by the Schrödinger wave function $f(x_{1},x_{2})\in{\mathcal{L}^{2}}(\mathbb{R}^{2},dx_{1}dx_{2})$ according to the definition [49, 50]
[…]
@@ 3.1 Wigner representation
where $p_{1}$ and $p_{2}$ are momentum variables, and $x_{1}$ and $x_{2}$ are position variables. The transfer of our original problem to the phase-space representation and the application of the WDF is partially motivated by the form of the Schrödinger wave functions of bipartite states given by Eqs (27) and (28). Let us note that both these wave functions have the form of the Gaussian functions shifted to different points of the configurational space and with different phase factors. This observation allows us to apply the following de Gosson result for the generalized Gaussian functions of the Wigner distribution function [50, Corollary 113] adapted to our needs, namely
Corollary 3.1
@@
$W(\psi_{\bf{M}})(\boldsymbol{\gamma})=\left(\frac{1}{\pi\hbar}\right)^{n}\exp{\left(-\frac{1}{2}{\boldsymbol{\gamma}}^{T}{\boldsymbol{\Sigma}^{-1}}{\boldsymbol{\gamma}}\right)},$ (33)
where ${{\boldsymbol{\Sigma}}}$ is the covariance matrix in the form
${\boldsymbol{\Sigma}}=\left[\begin{array}[]{ccc}\frac{1}{2}{\bf{M}}^{-1}&\bf{0}\\ \bf{0}&\frac{\hbar^{2}}{2}\bf{M}\end{array}\right],$ (34)
and $\boldsymbol{\gamma}^{T}$ is a $1\times n$ row matrix of the phase-space variables $(x_{1},x_{2},\ldots,x_{n},p_{1},p_{2},\ldots,p_{n})$.
@@
$\widetilde{\psi}_{k,\rm{coh}}^{(\alpha)}(x_{1},x_{2};z_{1},z_{2})=\widehat{T}({y}_{1,k},y_{2,k},q_{1,k},q_{2,k})\widetilde{\psi}_{k,\rm{coh}}^{(\alpha)}(x_{1},x_{2};0,0).$ (35)
According to Definition 2 in Ref. [50], the explicit form of the Heisenberg-Weyl operator is the following
@@
Let us note that the form of the wave functions given by Eqs (39) and (42) are consistent with the assumptions of the Corollary 3.1. Hence, the Wigner distribution functions corresponding to these wave functions can be directly expressed by Eq. (33), wherein the relations between matrices ${\bf{M}}_{k,\alpha}$ and ${\bf{\boldsymbol{\Sigma}}}_{k,\alpha}$ are given by Eq. (34), for $k=1,2$. However, the covariance matrices corresponding to these two cases are different. For the Wigner distribution function corresponding to the wave function $\widetilde{\psi}_{1,\rm{coh}}^{(\alpha)}(x_{1},x_{2};0,0)$, the covariance matrix has the form
${\bf{\boldsymbol{\Sigma}}}_{1,\alpha}=\frac{1}{2}\textrm{diag}\left(\frac{\alpha}{a^{2}},\frac{\alpha}{b^{2}},\frac{a^{2}\hbar^{2}}{\alpha},\frac{b^{2}\hbar^{2}}{\alpha}\right),$ (44)
@@
${\bf{\boldsymbol{\Sigma}}}_{2,\alpha}=\frac{1}{4\alpha}\left[\begin{array}[]{cccc}\frac{1+\alpha^{2}}{a^{2}}&\frac{\alpha^{2}-1}{ab}&0&0\\ \frac{\alpha^{2}-1}{ab}&\frac{1+\alpha^{2}}{b^{2}}&0&0\\ 0&0&(1+\alpha^{2})a^{2}\hbar^{2}&(1-\alpha^{2})ab\hbar^{2}\\ 0&0&(1-\alpha^{2})ab\hbar^{2}&(1+\alpha^{2})b^{2}\hbar^{2}\end{array}\right].$ (45)
Each of the obtained covariance matrices, ${\bf{\boldsymbol{\Sigma}}}_{i,\alpha}$ corresponding Gaussian’s Wigner distribution functions in the form given by Eq. (33), for $i=1,2$, satisfies the Robertson-Schrödinger uncertainty relation in the form
${\bf{\boldsymbol{\Sigma}}}_{i,\alpha}+\frac{i\hbar}{2}{\bf{J}}\geqslant 0,$ (46)
where ${\bf{J}}$ is the standard symplectic matrix (${\bf{J}}^{T}=-{\bf{J}}$ and ${\bf{J}}^{2}=-{\bf{I}}$, where ${\bf{I}}$ is the identity matrix). Let us note that the appearance of the matrix ${\bf{J}}$ in the aforementioned inequality is a consequence of the following position and momentum operators ordering in the row: $\hat{\boldsymbol{\gamma}}^{T}=[\hat{x}_{1}\;\hat{x}_{2}\;\hat{p}_{1}\;\hat{p}_{2}]$ and consequently, the canonical commutation relation takes the form: $\left[\hat{\gamma}_{k},\hat{\gamma}_{l}\right]=i\hbar J_{kl}{\textrm{Id}}_{L^{2}(\mathbb{R}^{2})}$, where $\hat{\gamma}_{k}$ refers to the components of vector $\hat{\boldsymbol{\gamma}}$, and $J_{kl}$ is the matrix element of the skew-symmetric matrix ${\bf{J}}$.
@@
${\widetilde{\bf{\boldsymbol{\Sigma}}}}_{i,\alpha}+\frac{i\hbar}{2}{\bf{J}}\geqslant 0,$ (47)
where the matrix ${\widetilde{{\boldsymbol{\Sigma}}}}_{i,\alpha}$ is given by the formula
${\widetilde{\bf{\boldsymbol{\Sigma}}}}_{i,\alpha}={\boldsymbol{\Lambda}}{{\bf{\boldsymbol{\Sigma}}}}_{i,\alpha}{\boldsymbol{\Lambda}}^{T}.$ (48)
In this formula, the matrix ${\boldsymbol{\Lambda}}$ has the form: ${\boldsymbol{\Lambda}}^{T}=\textrm{diag}\left(1,1,1,-1\right)$. On the other hand, de Gosson shows that the Simon-Horodecki-Peres separability criterion is equivalent to the following proposition [37, Proposition 168]
@@
where $\lambda_{min}$ is the minimal symplectic eigenvalue of the matrices ${\widetilde{\bf{\boldsymbol{\Sigma}}}_{i,\alpha}}$. Based on this proposition, we determine the symplectic eigenvalues of the matrices ${\widetilde{\boldsymbol{\Sigma}}}_{i,\alpha}$ by solving characteristic equations of the form
$\textrm{det}({\bf{J}}{\widetilde{\boldsymbol{\Sigma}}}_{i,\alpha}-\lambda{\bf{I}})=0.$ (50)
@@
However, let us note the eigenvalues of the matrix ${\bf{J}}{\widetilde{\boldsymbol{\Sigma}}}_{i,\alpha}$ are complex numbers in the form ±$\pm i\lambda^{(k)}_{i,\alpha}$ with $\lambda^{(k)}_{i,\alpha}>0$ , where the upper index $k$ numbers these eigenvalues. The positive numbers $\lambda^{(k)}_{i,\alpha}$ are the symplectic eigenvalues of the matrix ${\widetilde{\boldsymbol{\Sigma}}}_{i,\alpha}$ and they form the symplectic spectrum of this matrix. As a result of our calculations, we obtained the following results. For matrix ${\widetilde{\bf{\boldsymbol{\Sigma}}}_{1,\alpha}}$, we have one distinguishable symplectic eigenvalue of the form $\lambda^{(1)}_{1,\alpha}=\hbar/2$. In turn, for the matrix ${\widetilde{\bf{\boldsymbol{\Sigma}}}_{2,\alpha}}$ we have two different symplectic eigenvalues, namely $\lambda^{(1)}_{2,\alpha}=\hbar/(2\alpha)$ and $\lambda^{(2)}_{2,\alpha}=\hbar\alpha/2$. Based on these results, we can conclude that the symplectic eigenvalue $\lambda^{(1)}_{1,\alpha}=\hbar/2$ is equal to the minimal symplectic eigenvalue of the matrix ${\widetilde{\bf{\boldsymbol{\Sigma}}}_{1,\alpha}}$. Thus, the corresponding bipartite state has a separable form. However, this result was expected because the covariance matrix associated with this state is diagonal. On the other hand, both the symplectic eigenvalues of the matrix ${\widetilde{\bf{\boldsymbol{\Sigma}}}_{2,\alpha}}$ depend on the parameter $\alpha\in(0,1)$. Hence, the minimal eigenvalue of this matrix is $\lambda^{(min)}_{2,\alpha}=\lambda^{(2)}_{2,\alpha}$, and thereby the Simon-Horodecki-Peres separability criterion is violated. The bipartite state corresponding to this case is entangled. It is worth giving an in-depth at this result because it depends on the value of the $\alpha$ parameter. In particular, its limiting cases are quite interesting. Namely, in the limit, $\alpha\rightarrow 1_{-}$, the covariance matrix ${\bf{\boldsymbol{\Sigma}}}_{2,\alpha}$ is diagonal. It means that the corresponding bipartite state becomes disentangled. Moreover, it is also worth noting that in the limit $\alpha\rightarrow 1_{-}$, the holomorphic Hermite functions given by Eq. (22) are simplified to the product of the monomials in the form. In turn, in the limit $\alpha\rightarrow 0_{+}$, the covariance matrix, ${\bf{\boldsymbol{\Sigma}}}_{2,\alpha}$, corresponding to the aforementioned bipartite state remains non-diagonal, meaning the state is still entangled. However, the influence of this limit on the holomorphic Hermite functions (22) was examined only for the case $z_{2}=\bar{z}_{1}$, and it was shown that these functions are reduced to the form of the Ito polynomials [53].
@@
The application of the Simon-Horodecki-Peres criterion to the considered bipartite states enables us to identify one them as entangled, i.e. this one characterized by the covariance matrix ${\bf{\boldsymbol{\Sigma}}}_{2,\alpha}$. Unfortunately, it does not provide us a measure of their entanglement degree. To address this issue, we employ the logarithmic negativity measure [54, 55] in the form adapted to the bipartite Gaussian state, namely [56]
@@
3.2 Physical model realization

### v3

# Segal-Bargmann type spaces related to non-rotational measure, and entanglement of bipartite squeezed coherent states

@@ I Introduction
The presented description summarizes the main results of the one-mode picture of the squeezed coherent state construction. Applications and the role played by coherent states in quantum optics are described in Ref. [12, 13]. Other applications of the coherent states cover wide range of mathematical physics issues, starting from quantization [2, 14], through cosmology [15, 16], to field theory [17].
Another look at the coherence states and their properties comes from the multi-mode perspective. However, this exciting problem states a different thread we do not develop here because it is out of the scope of the proposed research. So, instead of this, we focus on the simplest case; namely, we consider the two-mode problem. It is admittedly more straightforward, but still offers the possibility of a deeper insight into the entanglement phenomena and the role of quantum coherence. Moreover, in some cases, the two-mode system seems to be an intriguing research playground for testing the concepts mentioned above. Recently, this issue has been examined regarding the factorization of bipartite Gaussian wave functions through the canonical transformation of position and momentum coordinates [18]. Admittedly, the authors demonstrated that this transformation can be used to factorize the wave functions under investigation. Still, their conclusions clearly show that the problem of entanglement of the bipartite states corresponding to these wave functions remains unsolved, because the mentioned transformation mixes degrees of freedom of the input subsystems on which they are originally defined. These results are generally consistent with the findings of Thirring et al. [19] and de la Torre et al. [20], that a problem of state separation or their entanglement depends on how observables are assigned to subsystems and how physical subsystems are identified. In other words, the discussed disentanglement of states resulting from the factorization of the families of the bipartite Gaussian functions in the previous work [18] is rather formal because the unitary transformation generated by the point transformation does not preserve the decomposition into subspaces associated with the input subsystems.
The structure of the bipartite states realized with the help of the squeezed coherent states opens a way to study the quantum states, containing entanglement and coherence on an equal footing. In other words, these two phenomena coexist in such a state through its construction.
@@
The novelty of the presented approach is that it examines the squeezing and entanglement phenomena starting from the description of bipartite states in the Bargmann space. The first step enables us to establish a direct connection between the parameter, $\alpha$, associated with the non-rotational measure in the Bargmann space and the squeezing parameter. By crossing with the calculations to the phase space and applying the results of Simon and de Gosson, we associate the phenomena of squeezing and entanglement. Moreover, we quantify the degree of entanglement using the log-negativity criterion using the symplectic eigenvalues [39] of the covariance matrix modified by Simon [40]. Our result explicitly depends on the parameter $\alpha$, thereby affirming the connection between the parameter associated with the non-rotational measure in the Bargmann space and entanglement. Finally, we extend the model presented in the previous work [18] to the case of two coupled quantum-mechanical anisotropic harmonic oscillators, with the coupling and anisotropy originating directly from the squeezing parameter. This result complements the study of applying coupled harmonic oscillators to investigate the problem between entanglement and squeezing in continuous-variable Gaussian systems [41]. Concurrently, it refers to the examination of the two coupled harmonic oscillators in the Segal-Bargmann space representation presented in Ref. [42].
The rest of this paper is organized as follows. In Sec. II, we recall selected properties of the holomorphic Hermite functions and discuss their utility in constructing single- and bipartite coherent and squeezed states. The main results of our study and their discussion are presented in Sec. III. In the first part of this section, we use the phase-space approach based on the Wigner distribution function to quantitatively investigate the entanglement issue in continuous-variable systems. In turn, in the second part, we develop a theoretical model of a two-particle mechanical system illustrating details of our study. Section IV contains a summary and conclusions. Finally, the paper concludes with an appendix containing the proofs of the Mehler formulae.
II Mathematical preliminaries
@@
$\frac{1}{\pi^{2}}\int_{\mathbb{C}^{2}}{dz_{1}dz_{2}}\;\overline{\widetilde{\psi}_{k,\rm{coh}}(w_{1}^{\prime},w_{2}^{\prime};z_{1},z_{2},\xi)}\,\widetilde{\psi}_{k,\rm{coh}}(w_{1},w_{2};z_{1},z_{2},\xi)=K(\bar{w}_{1},\bar{w}_{2},w^{\prime}_{1},w^{\prime}_{2}),$ (13)
@@
$\int_{\mathbb{C}}H_{m}(z)\overline{H_{n}(z)}e^{-(1-\alpha)x^{2}-(\frac{1}{\alpha}-1)y^{2}}dxdy=\frac{\pi\sqrt{\alpha}}{1-\alpha}\left(2\frac{1+\alpha}{1-\alpha}\right)^{n}n!\delta_{mn}.$ (20)
@@
The second class of analytic functions, $\varphi_{2,(m,n)}^{(\alpha)}(z_{1},z_{2})$, to be considered by us is generated by the holomorphic Hermite functions of the two complex variables, $h_{m,n}^{(\alpha)}(z_{1},z_{2})$, with $m,n=0,1,\ldots$, and $0<\alpha<1$, i.e. we assume that
@@
whose orthogonality is treated by [36, Theorem 3]. We emphasize that the Hermite functions (22) form the orthonormal basis with respect to the Gaussian measure (12). The algebraic and analytic properties of these polynomials are considered in Refs. [36, 45, 46, 47, 48]. In particular, it is also worth noting that the complex Hermite polynomials, $H_{m,n}(z_{1},z_{2})$ for $z_{2}=\bar{z}_{1}$ reduce themselves to the polynomials of the real variables $x,y$ with complex coefficient [44, 49, 50]. For the first time, this kind of polynomials, i.e. $H_{m,n}(z_{1},{\bar{z}_{1}})$ appeared in Ito’s work [49] on applications of complex random variables in stochastic processes. Hence, they are called the Ito polynomials. Let us note that the polynomials $H_{m,n}(z_{1},{\bar{z}_{1}})$ were also introduced and discussed in the physical context [51]. Finally, note that both the functions given by Eq. (17) and Eq.(21) satisfy the condition in the form
[…]
@@
III Results and discussion
III.1 Wigner representation
The phase-space representation is often used to examine physical systems, such as electromagnetic modes in optical or microwave cavities, or mechanical modes of harmonic oscillators, where degrees of freedom constitute continuous variables. This approach, frequently applied in quantum optics, is intuitive and robust, facilitating the investigation of the system non-classical state features or their dynamics. According to this approach, the system states within the phase space are represented by non-classical distribution functions of conjugate variables defined on a real symplectic space $\mathbb{R}^{2}\times\mathbb{R}^{2}$ equipped with canonical symplectic form, relevant to these variables. Among numerous such functions, the Wigner distribution function (WDF) plays a pivotal role, providing a lucid phase-space representation of the quantum state and emphasising the cognitive aspect of the obtained results. Let us first start by introducing the WDF for a bipartite pure state $\left|f\right$> represented by the Schrödinger wave function $f(x_{1},x_{2})\in{\mathcal{L}^{2}}(\mathbb{R}^{2},dx_{1}dx_{2})$ according to the definition [52, 53]
$W\left(f\right)(x_{1},x_{2},p_{1},p_{2})$ $=$ $\frac{1}{(2\pi\hbar)^{2}}\int_{\mathbb{R}^{2}}\!\!{dX_{1}dX_{2}}\exp\left[-\frac{i}{\hbar}\left(p_{1}X_{1}+p_{2}X_{2}\right)\right]$ (31)
@@ III.1 Wigner representation
where $p_{1}$ and $p_{2}$ are momentum variables, and $x_{1}$ and $x_{2}$ are position variables. The transfer of our original problem to the phase-space representation and the application of the WDF is partially motivated by the form of the Schrödinger wave functions of bipartite states given by Eqs (27) and (28). Let us note that both these wave functions have the form of the Gaussian functions shifted to different points of the configurational space and with different phase factors. This observation allows us to apply the following de Gosson result for the generalized Gaussian functions of the Wigner distribution function [53, Corollary 113] adapted to our needs, namely
Corollary III.1.
@@
$W(\psi_{\bf{M}})(\bm{\gamma})=\left(\frac{1}{\pi\hbar}\right)^{n}\exp{\left(-\frac{1}{2}{\bm{\gamma}}^{T}{\bm{\Sigma}^{-1}}{\bm{\gamma}}\right)},$ (33)
where $\bm{\Sigma}$}} is the covariance matrix in the form
${\bm{\Sigma}}=\left[\begin{array}[]{ccc}\frac{1}{2}{\bf{M}}^{-1}&\bf{0}\\ \bf{0}&\frac{\hbar^{2}}{2}\bf{M}\end{array}\right],$ (34)
and $\bm{\gamma}^{T}$ is a $1\times n$ row matrix of the phase-space variables $(x_{1},x_{2},\ldots,x_{n},p_{1},p_{2},\ldots,p_{n})$.
@@
$\widetilde{\psi}_{k,\rm{coh}}^{(\alpha)}(x_{1},x_{2};z_{1},z_{2})=\widehat{T}({y}_{1,k},y_{2,k},q_{1,k},q_{2,k})\widetilde{\psi}_{k,\rm{coh}}^{(\alpha)}(x_{1},x_{2};0,0),$ (35)
where $y_{j,k}$ and $q_{j,k}$ are the real and imaginary parts of complex numbers $z_{j}$, respectively, for $j=1,2$ and $k$ fixed. According to Definition 2 in Ref. [53], the explicit form of the Heisenberg-Weyl operator is the following
@@
Let us note that the form of the wave functions given by Eqs (39) and (42) are consistent with the assumptions of the Corollary III.1. Hence, the Wigner distribution functions corresponding to these wave functions can be directly expressed by Eq. (33), wherein the relations between matrices ${\bf{M}}_{k,\alpha}$ and ${\bf{\bm{\Sigma}}}_{k,\alpha}$ are given by Eq. (34), for $k=1,2$. However, the covariance matrices corresponding to these two cases are different. For the Wigner distribution function corresponding to the wave function $\widetilde{\psi}_{1,\rm{coh}}^{(\alpha)}(x_{1},x_{2};0,0)$, the covariance matrix has the form
${\bf{\bm{\Sigma}}}_{1,\alpha}=\frac{1}{2}\textrm{diag}\left(\frac{\alpha}{a^{2}},\frac{\alpha}{b^{2}},\frac{a^{2}\hbar^{2}}{\alpha},\frac{b^{2}\hbar^{2}}{\alpha}\right),$ (44)
@@
${\bf{\bm{\Sigma}}}_{2,\alpha}=\frac{1}{4\alpha}\left[\begin{array}[]{cccc}\frac{1+\alpha^{2}}{a^{2}}&\frac{\alpha^{2}-1}{ab}&0&0\\ \frac{\alpha^{2}-1}{ab}&\frac{1+\alpha^{2}}{b^{2}}&0&0\\ 0&0&(1+\alpha^{2})a^{2}\hbar^{2}&(1-\alpha^{2})ab\hbar^{2}\\ 0&0&(1-\alpha^{2})ab\hbar^{2}&(1+\alpha^{2})b^{2}\hbar^{2}\end{array}\right].$ (45)
Each of the obtained covariance matrices, ${\bf{\bm{\Sigma}}}_{k,\alpha}$ corresponding to Gaussian’s Wigner distribution functions in the form given by Eq. (33), for $k=1,2$, satisfies the Robertson-Schrödinger uncertainty relation in the form
${\bf{\bm{\Sigma}}}_{k,\alpha}+\frac{i\hbar}{2}{\bf{J}}\geqslant 0,$ (46)
where $\bf{J}$} is the standard symplectic matrix (${\bf{J}}^{T}=-{\bf{J}}$, ${\bf{J}}^{2}=-{\bf{I}}$, where $\bf{I}$} is the identity matrix). Let us note that the appearance of the matrix $\bf{J}$} in the aforementioned inequality is a consequence of the following position and momentum operators ordering in the row: $\hat{\bm{\gamma}}^{T}=[\hat{x}_{1}\;\hat{x}_{2}\;\hat{p}_{1}\;\hat{p}_{2}]$ and consequently, the canonical commutation relation takes the form: $\left[\hat{\gamma}_{k},\hat{\gamma}_{l}\right]=i\hbar J_{kl}{\textrm{Id}}_{L^{2}(\mathbb{R}^{2})}$, where $\hat{\gamma}_{k}$ refers to the components of vector $\hat{\bm{\gamma}}$, and $J_{kl}$ is the matrix element of the skew-symmetric matrix $\bf{J}$}.
@@
${\widetilde{\bf{\bm{\Sigma}}}}_{k,\alpha}+\frac{i\hbar}{2}{\bf{J}}\geqslant 0,$ (47)
where the matrix ${\widetilde{{\bm{\Sigma}}}}_{k,\alpha}$ is given by the formula
${\widetilde{\bf{\bm{\Sigma}}}}_{k,\alpha}={\bm{\Lambda}}{{\bf{\bm{\Sigma}}}}_{k,\alpha}{\bm{\Lambda}}^{T}.$ (48)
In this formula, the matrix $\bm{\Lambda}$} has the form: ${\bm{\Lambda}}^{T}=\textrm{diag}\left(1,1,1,-1\right)$. On the other hand, de Gosson shows that the Simon-Horodecki-Peres separability criterion is equivalent to the following proposition [39, Proposition 168]
@@
where $\lambda_{min}$ is the minimal symplectic eigenvalue of the matrices $\widetilde{\bf{\bm{\Sigma}}}_{k,\alpha}$}. Based on this proposition, we determine the symplectic eigenvalues of the matrices ${\widetilde{\bm{\Sigma}}}_{k,\alpha}$ by solving characteristic equations of the form
$\textrm{det}({\bf{J}}{\widetilde{\bm{\Sigma}}}_{k,\alpha}-\lambda{\bf{I}})=0.$ (50)
@@
However, let us note the eigenvalues of the matrix ${\bf{J}}{\widetilde{\bm{\Sigma}}}_{k,\alpha}$ are complex numbers in the form $\pm i\lambda^{(j)}_{k,\alpha}$ with $\lambda^{(j)}_{k,\alpha}>0$ , where the upper index $j$ numbers these eigenvalues. The positive numbers $\lambda^{(j)}_{k,\alpha}$ are symplectic eigenvalues of the matrix ${\widetilde{\bm{\Sigma}}}_{k,\alpha}$ and they form the symplectic spectrum of this matrix. As a result of our calculations, we obtained the following results. For matrix $\widetilde{\bf{\bm{\Sigma}}}_{1,\alpha}$}, we have one distinguishable symplectic eigenvalue of the form $\lambda^{(1)}_{1,\alpha}=\hbar/2$. In turn, for the matrix $\widetilde{\bf{\bm{\Sigma}}}_{2,\alpha}$} we have two different symplectic eigenvalues, namely $\lambda^{(1)}_{2,\alpha}=\hbar/(2\alpha)$ and $\lambda^{(2)}_{2,\alpha}=\hbar\alpha/2$. Based on these results, we can conclude that the symplectic eigenvalue $\lambda^{(1)}_{1,\alpha}=\hbar/2$ is equal to the minimal symplectic eigenvalue of the matrix $\widetilde{\bf{\bm{\Sigma}}}_{1,\alpha}$}. Thus, the corresponding bipartite state has a separable form. However, this result was expected because the covariance matrix associated with this state is diagonal. On the other hand, both the symplectic eigenvalues of the matrix $\widetilde{\bf{\bm{\Sigma}}}_{2,\alpha}$} depend on the parameter $\alpha\in(0,1)$. Hence, the minimal eigenvalue of this matrix is $\lambda^{(min)}_{2,\alpha}=\lambda^{(2)}_{2,\alpha}$, and thereby the Simon-Horodecki-Peres separability criterion is violated. The bipartite state corresponding to this case is entangled. It is worth giving an in-depth look at this result because it depends on the value of the $\alpha$ parameter. In particular, its limiting cases are quite interesting. Namely, in the limit, $\alpha\rightarrow 1_{-}$, the covariance matrix ${\bf{\bm{\Sigma}}}_{2,\alpha}$ is diagonal. It means that the corresponding bipartite state becomes disentangled. Moreover, it is also worth noting that in the limit $\alpha\rightarrow 1_{-}$, the holomorphic Hermite functions given by Eq. (22) are simplified to the product form of the monomials. In turn, in the limit $\alpha\rightarrow 0_{+}$, the covariance matrix, ${\bf{\bm{\Sigma}}}_{2,\alpha}$, corresponding to the aforementioned bipartite state remains non-diagonal, meaning the state is still entangled. However, the influence of this limit on the holomorphic Hermite functions (22) was examined only for the case $z_{2}=\bar{z}_{1}$, and it was shown that these functions are reduced to the form of the Ito polynomials [51].
@@
The application of the Simon-Horodecki-Peres criterion to the considered bipartite states enables us to identify one them as entangled, i.e. this one characterized by the covariance matrix ${\bf{\bm{\Sigma}}}_{2,\alpha}$. Unfortunately, it does not provide us a measure of their entanglement degree. To address this issue, we employ the logarithmic negativity measure [56, 57] in the form adapted to the bipartite Gaussian state, namely [58]
@@
III.2 Physical model realization